### v1

@@ Abstract
Methods. In this work, we use NLTE line formation, as well as one-dimensional (1D) hydrostatic and spatial averages of 3D hydrodynamical (3D) model atmospheres to measure the abundance of Eu based on both the Eu II 4129 Å and Eu II 6645 Å lines for solar spectra and metal-poor stars.
Results. We find that for Eu II 4129 Å line the NLTE modelling leads to higher (0.04 dex) solar Eu abundance in 1D and higher (0.07 dex) in 3D NLTE while NLTE modelling leads to higher (0.01 dex) solar Eu abundance in 1D and lower (0.03 dex) in 3D NLTE for Eu II 6645 Å line. Although the NLTE corrections for the Eu II 4129 Å and Eu II 6645 Å lines are opposite, the discrepancy between the abundances derived from these individual lines reduces after applying NLTE corrections, highlighting the critical role of NLTE abundance determinations. By comparing these measurements with Galactic chemical evolution (GCE) models, we find that the amount of NLTE correction does not require significant change of the parameters for Eu production in the GCE models.
@@
Abundance ratios as a function of time or metallicity provide important information about the scenarios of nucleosynthesis and also allow us to trace the chronology of events in Galaxy, since different chemical elements are produced on different time scales in different astrophysical sites (Rana 1991; Wilson & Rood 1994; Chiappini et al. 2003; Matteucci 2012, 2014). Iron peak elements such as iron, cobalt, and chromium, are produced over longer time-scales in Type Ia supernovae (Burbidge et al. 1957; Nomoto et al. 1994; Hillebrandt & Niemeyer 2000; Seitenzahl et al. 2013; Bergemann et al. 2019a). In addition, the abundance of -elements such as magnesium and oxygen has been demonstrated to be a sensitive proxy for the initial phases of Galactic chemical evolution, which are dominated by the production of heavy elements and the return from core-collapse supernovae (Tinsley 1980; Woosley & Heger 2015). Slow neutron-capture process (s-process) elements, like strontium and barium, are produced in intermediate and low-mass asymptotic giant branch (AGB) stars (Tinsley 1980; Pagel 1997; Matteucci 2001; Karakas & Lattanzio 2014; Frischknecht et al. 2016; Choplin et al. 2018). Similar, rapid neutron-capture process (r-process) elements, such as europium, provide insight into violent events such as neutron star mergers, neutron-driven winds in core-collapse SNe, explosions of rapidly rotating magnetised massive stars (Rosswog et al. 1999; Takahashi et al. 1994; Woosley et al. 1994; Arcones & Thielemann 2013; Bliss et al. 2018; Siegel et al. 2019; Reichert et al. 2023).
@@
Guiglion et al. (2018) investigated three pure r-process elements, including Eu, Gd, and Dy based on MARCS model atmosphere and the local thermodynamic equilibrium (LTE) code. They found that the [Eu/Fe] ratio follows a continuous sequence from the thin disk to the thick disk with respect to metallicity. Lian et al. (2023) compared the evolution of the [Eu/Fe]-[Fe/H] trend, based on Eu abundance measurements using 1D LTE model, with the GCE model in the metal-rich region. However, their study lacks a metal-poor sample and relies solely on 1D LTE results. As shown in recent literature studies (Bergemann & Nordlander 2014a; Bergemann & Nordlander 2014b; Bergemann et al. 2017a; Storm & Bergemann 2023; Storm et al. 2024), it is crucial to take into account the effects of NLTE and 3D effects for FGKM-type stars. Therefore, further study of applying NLTE and 3D corrections to Eu abundances in a metal-poor sample is essential.
The first analysis of NLTE effects in Eu can be traced back to Mashonkina & Gehren (2000), which presents the statistical equilibrium of Eu II using a model atom that includes 32 levels of Eu II along with the ground state of Eu III and derives the solar europium abundance with A(Eu) = 0.53 dex. Zhao et al. (2016) follow the atomic model from Mashonkina & Gehren (2000) and investigate Eu abundances in 1D NLTE analysis. Recently, Storm et al. (2024) published the latest Eu atom model with 163 levels of Eu II to investigate the solar abundance of Eu. In this work, we apply NLTE line formation, as well as 1D hydrostatic and 3D model atmospheres with the latest Eu atomic data from Storm et al. (2024) to measure the abundance of Eu of 164 metal-poor stars in the Galactic halo and metal-poor disk with [Fe/H] ranging from to dex adopted from Ruchti et al. (2011).
@@ 2 Data
The sample is adopted from Ruchti et al. (2011). It includes 319 metal-poor stars in Galactic disk and halo with the effective temperature () ranging from to K, the surface gravity () ranging from to , the metallicities ranging from to dex. These stars were selected to investigate thick disk-like kinematics, and were originally observed by the RAVE survey (Steinmetz et al. 2006). The sample was observed using several high-resolution facilities, include the MIKE spectrograph on the Magellan-Clay telescope (Bernstein et al. 2003), FEROS on the MPG 2.2 m telescope (Kaufer et al. 1999), the UCLES spectrograph on the Anglo-Australian telescope (Walker & Diego 1985), and the ARC spectrograph on the Apache Point 3.5 m telescope (Wang et al. 2003). The wavelength range of the spectra observed from UCLES is from 4460 to 7270 Å, while the other three cover the full optical range from 3500 to 9500 Å. The signal-to-noise ratios (S/N) of all the spectra are greater than 100 pixel-1 at around 6000 Å. We adopted the NLTE-opt stellar parameters from Table 1 in Bergemann et al. (2017b) as the input data for our further analysis.
@@ 3 Methods
Here, we will provide a description of the model atmospheres, spectral synthesis code, NLTE models, and linelist that we use in this work.
@@ 3.1 Model Atmospheres
The MARCS model is a homogeneous model atmosphere for late-type (FGKM) stars, assuming LTE. This model uses the thermal equilibrium and Saha equation for all number densities of atoms and molecules (Gibson & Heitler 1928; Russell 1934), as well as the Boltzmann distribution for all partition functions and excitation equilibrium (Irwin 1981; Gustafsson et al. 2008). The MARCS model atmospheres provided standard composition model atmospheres111https://marcs.astro.uu.se covering a wide parameter space: The range of effective temperature () are from 2500K to 8000K, with a step of 100K for () from 2500K to 4000K and a step of 250K for all other ones. For surface gravity (log g), the range is from to dex with a step of dex. The range of metallicity ([Fe/H]) is from to , with a step of dex from to dex, dex from to , and dex from to . The micro-turbulence, with four different values of , , and km , is also included in the grids.
@@ 3.2 Spectral synthesis code
A Python wrapper called Turbospectrum Spectral Fitting with Python (TSFitPy)444https://github.com/JGerbs13/TSFitPy developed by Gerber et al. (2023) is specifically designed to determine stellar abundances, optional with other parameters such as micro-turbulence () using the Nelder-Mead (simplex algorithm) minimization method (Nelder & Mead 1965; Wright 1996). TSFitPy was notably updated to fit either macroturbulence or rotation for each individual line using L-BFGS-B (Limited-memory Broyden-Fletcher-Goldfarb-Shanno with bound consideration) algorithm (Byrd et al. 1995; Zhu et al. 1997) using the Scipy Python package (Virtanen et al. 2020). This is done as a secondary step after generating synthetic spectra with a specific stellar abundance to break the degeneracy of fitting the abundance and the broadening of lines simultaneously (Storm & Bergemann 2023). A dedicated interpolating function is provided together with TS code, which takes the rectangular grids of model atmospheres and the corresponding grids of departure coefficients, and produces an interpolated atmosphere structure and departure file for a desired combination of stellar parameters (Gerber et al. 2023).
In this work, we use the updated TSFitPy and TS code to fit observed spectra and obtain the stellar abundances of the metal-poor sample.
@@ 3.3 NLTE models
We use MULTI1D code to compute statistical equilibrium (SE) calculations, which was developed by Carlsson (1986). MULTI1D code solves the the SE equations iteratively and handles the radiative transfer equation in a 1D plane-parallel geometry, assuming that deviations from LTE do not influence the structure of the input model atmosphere, which underlies the standard assumption of a trace element to calculate the statistical equilibrium of NLTE elements. Recently, the MULTI1D was updated by our group (Bergemann et al. 2019a; Gallagher et al. 2020) and are widely used for NLTE analyses of atmospheric parameters and chemical abundances (Gerber et al. 2023; Storm & Bergemann 2023; Li & Ezzeddine 2023).
We use a Python wrapper777https://github.com/stormnick/wrapper_multi for MULTI1D to calculate the departure coefficients of Eu for both grids of MARCS and 3D Stagger model atmospheres (using the same methodology as in Sec. 2.4 in Gerber et al. 2023). For each individual star, we obtain the abundance directly using TS by fitting the spectra, where the NLTE synthetic spectra are based on the precomputed departure coefficients from MULTI1D. As detailed, the europium departure coefficient grids are computed between [Eu/Fe] = to +1 in steps of 0.1 dex. During the fitting process, the closest departure coefficient within [Eu/Fe] step is used ( within 0.05 dex) for the synthetic spectra generation. For example a star with [Eu/Fe] = 0.6 dex would use departure coefficients computed with that specific abundance. Therefore, different [Eu/Fe] abundances are taken into account during the statistical equilibrium calculations for each individual star.
@@ 3.4 Linelist and line choice
We use the homogeneous Gaia-ESO linelist, which contains atomic and molecular data from Heiter et al. (2021) and was recently updated by Magg et al. (2022) with atomic data for several elements. For the solar Eu abundance, we adopt A(Eu) = 0.57 from Storm et al. (2024). For Eu II lines, the -values were measured by Lawler et al. (2001) based on experimental life-times and branching fractions (BFs).
@@
There are two odd isotopes for Eu which are and . The solar isotopic abundance ratio for this two isotopes is set to 47.8:52.2, respectively from Lodders et al. (2009). The Eu II 4129Å line is represented by 11 individual hyperfine splitting (HFS) components from Mashonkina & Gehren (2000). The Eu II 6645Å line is represented by 11 hyperfine and isotopic components (Storm et al. 2024).
@@ 4.1 Synthetic spectra
To showcase TSFitPy capability, we generated the spectra with 1D LTE/NLTE and 3D LTE/NLTE lines profiles based on solar parameters for Eu II 4129Å (left) and Eu II 6645Å (right) in Figure 2. The red solid lines represent line profiles generated from 1D LTE, while the blue solid lines are from 1D NLTE. The red dashed lines represent line profiles generated from 3D LTE, and the blue dashed lines are from 3D NLTE. From the comparison of the red and blue lines in the left panel, we observe that the NLTE effect weakens the Eu II 4129Å line in both 1D and 3D models. By comparing the red and blue solid lines in the right panel, we find that the NLTE effect slightly weakens the line in the 1D model. This result is consistent with the positive 1D NLTE correction found by Mashonkina & Gehren (2000) and Storm et al. (2024). In contrast, when comparing the red and blue dashed lines, we observe that the NLTE effect strengthens the line in the 3D model. Storm et al. (2024) studied the impact of full 3D NLTE modeling of Eu on solar abundances. They found that the Eu NLTE model results in a slightly positive correction in 1D and a negative correction in 3D, which is consistent with our findings.
@@ 4.2 Solar Eu abundance
In the top panels of Fig. 3, we show the observed Sun spectrum of the selected Eu II 4129Å line (left) and Eu II 6645Å (right) line in black dots, taken from the Vacuum Vertical Telescope at the Institute für Astrophysik Göttingen (IAG) with the resolving power of R = 700,000 (Reiners et al. 2016). We overplot the synthetic spectrum based on the 1D NLTE model with the best-fit A(Eu) = 0.61 for 4129Å line and A(Eu) = 0.55 for 6645Å line in red lines, and the synthetic spectrum based on the 3D NLTE model with the best-fit A(Eu) = 0.64 for 4129Å line and A(Eu) = 0.58 for 6645Å line in blue lines. We accounted for the blending of the Eu II 4129.7 Å line by Ti I 4129.66 Å line and Sc I 4129.75 (Mashonkina & Gehren 2000). For Eu II 6645.70 Å line, we accounted for the blending of Si I 6645.21 line. For different stars in Sec. 4.4, the Ti, Sc and Si abundance was scaled to [Fe/H]. The individual error components were calculated by adding the systematic uncertainty (including blends and uncertainty of -values) and the fitting error in quadrature (Storm et al. 2024). In the bottom panels of Fig. 3, we present the observed Sun spectrum in black dots, derived from the solar Kitt Peak National Observatory (KPNO) FTS atlas with R 400,000 (Kurucz et al. 1984). The synthetic spectrum based on the 1D NLTE model with the best-fit A(Eu) = 0.62 for 4129Å line A(Eu) = 0.58 for 6645Å line are overplotted in red, and the synthetic spectrum based on the 3D NLTE model with the best-fit A(Eu) = 0.64 for 4129Å line and A(Eu) = 0.60 for 6645Å line are shown in blue. This indicates that the Eu abundances derived from the 1D NLTE model are consistent across the two observed spectra, and similarly, the results from the 3D NLTE model are also consistent. The best-fit values of 1D LTE and 3D LTE, are also provided in Table. 2.
@@
Our results for 1D LTE and NLTE abundances are higher than those reported by Mashonkina & Gehren (2000), who obtained LTE and NLTE abundances for 4129Å line are 0.49 and 0.53 and for 6645Å line are 0.50 and 0.53, respectively. The differences could be attributed to several factors, including the different loggf values used in our study (0.22 for 4129Å line and 0.12 for 6645Å line from Lawler et al. (2001); see Section 3.3), compared to the value of 0.174 for 4129Å and 0.204 for 6645Å taken from Komarovskii (1991) in Mashonkina & Gehren (2000). Additionally, the Eu II lines are also influenced by hyperfine structure and isotopic shifts. We used hyperfine structure data described in Section 3.4 for the Eu II 6645Å line, represented with 11 hyperfine and isotopic components from Storm et al. (2024), whereas Mashonkina & Gehren (2000) used data from Biehl (1976). The Eu isotope ratio used in Mashonkina & Gehren (2000) was 55:45, whereas we used a ratio of 47.8:52.2 from Lodders et al. (2009) based on the Gaia-ESO line list (Heiter et al. 2021). Furthermore, we used the MULTI1D code to calculate the statistical equilibrium of Eu, while Mashonkina & Gehren (2000) used the NONLTE3 code. Besides, due to these differences in atomic data and the statistical equilibrium code, Storm et al. (2024) tested the main factors influencing the results and concluded that they are primarily due to the use of a more comprehensive atomic model.
@@
The difference for Eu II 6645Å line in abundance corrections is consistent with findings from Storm et al. (2024), showing an increasing between 1D LTE and 1D NLTE, and a decreasing from 3D (full 3D) LTE to 3D (full 3D) NLTE. However, our 3D results are higher than their full 3D results, which are 0.58 and 0.55. This may indicate that 3D cannot represent all of the 3D effects, but nevertheless provides a great insight into its effects.

### v2

@@ Abstract
Methods. In this work, we used NLTE line formation, as well as one-dimensional (1D) hydrostatic and spatial averages of three-dimensional hydrodynamical (3D) model atmospheres to measure the abundance of Eu based on both the Eu II 4129 Å and Eu II 6645 Å lines for solar spectra and metal-poor stars.
Results. We find that for Eu II 4129 Å line the NLTE modeling leads to higher (0.04 dex) solar Eu abundance in 1D and higher (0.07 dex) in 3D NLTE while NLTE modeling leads to higher (0.01 dex) solar Eu abundance in 1D and lower (0.03 dex) in 3D NLTE for Eu II 6645 Å line. Although the NLTE corrections for the Eu II 4129 Å and Eu II 6645 Å lines are opposite, the discrepancy between the abundances derived from these individual lines reduces after applying NLTE corrections, highlighting the critical role of NLTE abundance determinations. By comparing these measurements with Galactic chemical evolution (GCE) models, we find that the amount of NLTE correction does not require significant change of the parameters for Eu production in GCE models.
@@
Abundance ratios as a function of time or metallicity provide important information about the scenarios of nucleosynthesis and also allow us to trace the chronology of events in Galaxy, since different chemical elements are produced on different time scales in different astrophysical sites (Rana 1991; Wilson & Rood 1994; Chiappini et al. 2003; Matteucci 2012, 2014). Iron peak elements such as iron, cobalt, and chromium, are produced over longer time-scales in Type Ia supernovae (Burbidge et al. 1957; Nomoto et al. 1994; Hillebrandt & Niemeyer 2000; Seitenzahl et al. 2013; Bergemann et al. 2019a). In addition, the abundance of -elements such as magnesium and oxygen has been demonstrated to be a sensitive proxy for the initial phases of Galactic chemical evolution, which are dominated by the production of heavy elements and the return from core-collapse supernovae (Tinsley 1980; Woosley & Heger 2015). Slow neutron-capture process (s-process) elements, such as strontium and barium, are produced in intermediate and low-mass asymptotic giant branch (AGB) stars (Tinsley 1980; Pagel 1997; Matteucci 2001; Karakas & Lattanzio 2014; Frischknecht et al. 2016; Choplin et al. 2018). Similar, rapid neutron-capture process (r-process) elements, such as europium, provide insight into violent events such as neutron star mergers, neutron-driven winds in core-collapse SNe, explosions of rapidly rotating magnetised massive stars (Rosswog et al. 1999; Takahashi et al. 1994; Woosley et al. 1994; Arcones & Thielemann 2013; Bliss et al. 2018; Siegel et al. 2019; Reichert et al. 2023).
@@
Guiglion et al. (2018) investigated three pure r-process elements, including Eu, Gd, and Dy based on MARCS model atmosphere and the local thermodynamic equilibrium (LTE) code. They found that the [Eu/Fe] ratio follows a continuous sequence from the thin disk to the thick disk with respect to metallicity. Lian et al. (2023) compared the evolution of the [Eu/Fe]-[Fe/H] trend, based on Eu abundance measurements using 1D LTE model, with the GCE model in the metal-rich region. However, their study lacks a metal-poor sample and relies solely on 1D LTE results. As shown in recent literature studies (Bergemann & Nordlander 2014a; Bergemann & Nordlander 2014b; Bergemann et al. 2017a; Storm & Bergemann 2023; Storm et al. 2024), it is crucial to take into account the effects of NLTE and 3D effects for FGKM-type stars. Therefore, further studies based on the application of NLTE and 3D corrections to Eu abundances in a metal-poor sample is essential.
The first analysis of NLTE effects in Eu can be traced back to Mashonkina & Gehren (2000), which presents the statistical equilibrium of Eu II using a model atom that includes 32 levels of Eu II, along with the ground state of Eu III. This approach derived the solar europium abundance with A(Eu) = 0.53 dex. Zhao et al. (2016) followed the atomic model from Mashonkina & Gehren (2000) and investigated Eu abundances in 1D NLTE analysis. Recently, Storm et al. (2024) published the latest Eu atom model with 163 levels of Eu II to investigate the solar abundance of Eu. In this work, we apply NLTE line formation, as well as 1D hydrostatic and 3D model atmospheres with the latest Eu atomic data from Storm et al. (2024) to measure the abundance of Eu of 164 metal-poor stars in the Galactic halo and metal-poor disk with [Fe/H] ranging from to dex adopted from Ruchti et al. (2011).
@@ 2 Data
The sample is adopted from Ruchti et al. (2011). It includes 319 metal-poor stars in Galactic disk and halo with the effective temperature () ranging from to K, the surface gravity () ranging from to , the metallicities ranging from to dex. These stars were selected to investigate thick disk-like kinematics, and they were originally observed by the RAVE survey (Steinmetz et al. 2006). The sample was observed using several high-resolution facilities, including the MIKE spectrograph on the Magellan-Clay telescope (Bernstein et al. 2003), FEROS on the MPG 2.2 m telescope (Kaufer et al. 1999), UCLES spectrograph on the Anglo-Australian telescope (Walker & Diego 1985), and the ARC spectrograph on the Apache Point 3.5 m telescope (Wang et al. 2003). The wavelength range of the spectra observed from UCLES is from 4460 to 7270 Å, whereas the other three cover the full optical range from 3500 to 9500 Å. The signal-to-noise ratios (S/Ns) of all the spectra are greater than 100 pixel-1 at around 6000 Å. We adopted the NLTE-opt stellar parameters from Table 1 in Bergemann et al. (2017b) as the input data for further analysis.
@@ 3 Methods
Here, we provide a description of the model atmospheres, spectral synthesis code, NLTE models, and the linelist we use in this work.
@@ 3.1 Model atmospheres
The MARCS model is a homogeneous model atmosphere for late-type (FGKM) stars, assuming LTE. This model uses the thermal equilibrium and Saha equation for all number densities of atoms and molecules (Gibson & Heitler 1928; Russell 1934), as well as the Boltzmann distribution for all partition functions and excitation equilibrium (Irwin 1981; Gustafsson et al. 2008). The MARCS model atmospheres provided standard composition model atmospheres111https://marcs.astro.uu.se covering a wide parameter space: the range of effective temperature () goes from 2500K to 8000K, with a step of 100K for () from 2500K to 4000K and a step of 250K for all others. For the surface gravity (log g), the range is from to dex with a step of dex. The range of metallicity ([Fe/H]) is from to , with a step of dex from to dex, dex from to , and dex from to . The micro-turbulence, with four different values (, , , and km ) is also included in the grids.
@@ 3.2 Spectral synthesis code
A Python wrapper called Turbospectrum Spectral Fitting with Python (TSFitPy)444https://github.com/JGerbs13/TSFitPy developed by Gerber et al. (2023) has specifically been designed to determine stellar abundances, optional with other parameters, such as micro-turbulence () using the Nelder-Mead (simplex algorithm) minimization method (Nelder & Mead 1965; Wright 1996). TSFitPy was notably updated to fit either macroturbulence or rotation for each individual line using Limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS-B) with bound consideration algorithm (Byrd et al. 1995; Zhu et al. 1997) using the Scipy Python package (Virtanen et al. 2020). This was done as a secondary step after generating synthetic spectra with a specific stellar abundance to break the degeneracy of fitting the abundance and the broadening of lines simultaneously (Storm & Bergemann 2023). A dedicated interpolating function is provided together with TS code, which takes the rectangular grids of model atmospheres and the corresponding grids of departure coefficients, and produces an interpolated atmosphere structure and departure file for a desired combination of stellar parameters (Gerber et al. 2023).
In this work, we used the updated TSFitPy and TS code to fit the observed spectra and obtain the stellar abundances of the metal-poor sample.
@@ 3.3 NLTE models
We used the MULTI1D code to compute statistical equilibrium (SE) calculations, which was developed by Carlsson (1986). This code solves the the SE equations iteratively and handles the radiative transfer equation in a 1D plane-parallel geometry, assuming that deviations from LTE do not influence the structure of the input model atmosphere, which underlies the standard assumption of a trace element to calculate the statistical equilibrium of NLTE elements. Recently, MULTI1D was updated by our group (Bergemann et al. 2019a; Gallagher et al. 2020) and widely used for NLTE analyses of atmospheric parameters and chemical abundances (Gerber et al. 2023; Storm & Bergemann 2023; Li & Ezzeddine 2023).
We used a Python wrapper777https://github.com/stormnick/wrapper_multi for MULTI1D to calculate the departure coefficients of Eu for both grids of MARCS and 3D Stagger model atmospheres (using the same methodology as in Sec. 2.4 in Gerber et al. 2023). For each individual star, we obtained the abundance directly using TS by fitting the spectra, where the NLTE synthetic spectra are based on the precomputed departure coefficients from MULTI1D. As detailed, the europium departure coefficient grids are computed between [Eu/Fe] = to +1 in steps of 0.1 dex. During the fitting process, the closest departure coefficient within [Eu/Fe] step is used ( within 0.05 dex) for the synthetic spectra generation. For example a star with [Eu/Fe] = 0.6 dex would use departure coefficients computed with that specific abundance. Therefore, different [Eu/Fe] abundances are taken into account during the statistical equilibrium calculations for each individual star.
@@ 3.4 Linelist and line choice
We used the homogeneous Gaia-ESO linelist, which contains atomic and molecular data from Heiter et al. (2021) and was recently updated by Magg et al. (2022) with atomic data for several elements. For the solar Eu abundance, we adopted A(Eu) = 0.57 from Storm et al. (2024). For the Eu II lines, the -values were measured by Lawler et al. (2001) based on experimental life-times and branching fractions (BFs).
@@
There are two odd isotopes for Eu, namely, and . The solar isotopic abundance ratio for this two isotopes is set to 47.8:52.2, respectively from Lodders et al. (2009). The Eu II 4129Å line is represented by 11 individual hyperfine splitting (HFS) components from Mashonkina & Gehren (2000). The Eu II 6645Å line is represented by 11 hyperfine and isotopic components (Storm et al. 2024).
@@ 4.1 Synthetic spectra
To showcase the TSFitPy capability, we generated the spectra with 1D LTE/NLTE and 3D LTE/NLTE lines profiles based on solar parameters for Eu II 4129Å (left) and Eu II 6645Å (right) in Figure 2. The red solid lines represent line profiles generated from 1D LTE, while the blue solid lines are from 1D NLTE. The red dashed lines represent line profiles generated from 3D LTE, and the blue dashed lines are from 3D NLTE. From the comparison of the red and blue lines in the left panel, we observe that the NLTE effect weakens the Eu II 4129Å line in both 1D and 3D models. By comparing the red and blue solid lines in the right panel, we find that the NLTE effect slightly weakens the line in the 1D model. This result is consistent with the positive 1D NLTE correction found by Mashonkina & Gehren (2000) and Storm et al. (2024). In contrast, when comparing the red and blue dashed lines, we observe that the NLTE effect strengthens the line in the 3D model. Storm et al. (2024) studied the impact of full 3D NLTE modeling of Eu on solar abundances. They found that the Eu NLTE model results in a slightly positive correction in 1D and a negative correction in 3D, which is consistent with our findings.
@@ 4.2 Solar Eu abundance
In the top panels of Fig. 3, we show the observed Sun spectrum of the selected Eu II 4129Å line (left) and Eu II 6645Å (right) line in black dots, taken from the Vacuum Vertical Telescope at the Institute für Astrophysik Göttingen (IAG), with a resolving power of R = 700,000 (Reiners et al. 2016). Here, we overplot the synthetic spectrum based on the 1D NLTE model with the best-fit A(Eu) = 0.61 for 4129Å line and A(Eu) = 0.55 for 6645Å line in red lines. The synthetic spectrum based on the 3D NLTE model with the best-fit A(Eu) = 0.64 for 4129Å line and A(Eu) = 0.58 for 6645Å line is shown as the blue lines. We accounted for the blending of the Eu II 4129.7 Å line by Ti I 4129.66 Å line and Sc I 4129.75 (Mashonkina & Gehren 2000). For Eu II 6645.70 Å line, we accounted for the blending of Si I 6645.21 line. For different stars in Sec. 4.4, the Ti, Sc and Si abundance was scaled to [Fe/H]. The individual error components were calculated by adding the systematic uncertainty (including blends and uncertainty of -values) and the fitting error in quadrature (Storm et al. 2024). In the bottom panels of Fig. 3, we present the observed Sun spectrum in black dots, derived from the solar Kitt Peak National Observatory (KPNO) FTS atlas with R 400,000 (Kurucz et al. 1984). The synthetic spectrum based on the 1D NLTE model with the best-fit A(Eu) = 0.62 for 4129Å line A(Eu) = 0.58 for the 6645Å line are overplotted in red, while the synthetic spectrum based on the 3D NLTE model with the best-fit A(Eu) = 0.64 for 4129Å line and A(Eu) = 0.60 for 6645Å line are shown in blue. This indicates that the Eu abundances derived from the 1D NLTE model are consistent across the two observed spectra, and similarly, the results from the 3D NLTE model are also consistent. The best-fit values of 1D LTE and 3D LTE, are also provided in Table. 2.
@@
Our results for 1D LTE and NLTE abundances are higher than those reported by Mashonkina & Gehren (2000), who obtained LTE and NLTE abundances for 4129Å line are 0.49 and 0.53, while for the 6645Å line, these are 0.50 and 0.53, respectively. The differences could be attributed to several factors, including the different loggf values used in our study (0.22 for 4129Å line and 0.12 for 6645Å line from Lawler et al. (2001); see Section 3.3), compared to the value of 0.174 for 4129Å and 0.204 for 6645Å taken from Komarovskii (1991) in Mashonkina & Gehren (2000). Additionally, the Eu II lines are also influenced by hyperfine structure and isotopic shifts. We used the hyperfine structure data described in Section 3.4 for the Eu II 6645Å line, represented with 11 hyperfine and isotopic components from Storm et al. (2024); whereas Mashonkina & Gehren (2000) used data from Biehl (1976). The Eu isotope ratio used in Mashonkina & Gehren (2000) was 55:45, whereas we used a ratio of 47.8:52.2 from Lodders et al. (2009) based on the Gaia-ESO line list (Heiter et al. 2021). Furthermore, we used the MULTI1D code to calculate the statistical equilibrium of Eu, while Mashonkina & Gehren (2000) used the NONLTE3 code. In addition, due to these differences in atomic data and the statistical equilibrium code, Storm et al. (2024) tested the main factors influencing the results and concluded that they are primarily due to the use of a more comprehensive atomic model.
@@
The difference for the Eu II 6645Å line in abundance corrections is consistent with findings from Storm et al. (2024), showing an increasing between 1D LTE and 1D NLTE, while decreasing from 3D (full 3D) LTE to 3D (full 3D) NLTE. However, our 3D results are higher than their full 3D results, which are 0.58 and 0.55. This may indicate that 3D cannot represent all of the 3D effects, but nevertheless provides a promising insight into its effects.